OpenStudy (anonymous):

if the sum of binomial coefficient in expansion of [xpower2/3-2/x]power n is 4096, then find greatest binomial coefficient and also 5th term.

OpenStudy (agent0smith):

\[\left( \frac{ x^2 }{ 3 }- \frac{ 2 }{x } \right)^{4096}\] is that what it looks like...?

OpenStudy (anonymous):

yup